# Dominik Bojko Krzysztof Grining Marek Klonowski

### v6

@@ Abstract
We demonstrate that probabilistic counters can be used as a privacy protection mechanism without extra randomization. Namely, the inherent randomization from the protocol is sufficient for protecting privacy, even if the probabilistic counter is used multiple times. In particular, we present a specific privacy-preserving data aggregation protocol based on Morris Counter and MaxGeo Counter. Some of the presented results are devoted to counters that have not been investigated so far from the perspective of privacy protection. Another part is an improvement of previous results. We show how our results can be used to perform distributed surveys and compare the properties of counter-based solutions and a standard Laplace method.
@@
differential privacy, data aggregation, probabilistic counters
@@
[pwr]Wroclaw University of Science and Technology, Faculty of Fundamental Problems of Technology, Department of Fundamentals of Computer Science
[…]
[pwr-ai]Wroclaw University of Science and Technology, Faculty of Fundamental Problems of Technology, Department of Artificial Intelligence
@@ 1 Introduction
Since Big Data related topics have been widely developed in recent years, solutions that focus on saving memory resources have become very popular. We would like to consider a standard example of such space-efficient mechanisms, namely probabilistic counters, which are used to represent the cardinality of dynamically counted events. More precisely, we would like to indicate the occurrence of $n$ events using a very small (significantly less than $\log n$) number of bits. We assume that $n$ is unknown in advance and may change. Clearly, a simple information-theoretic argument convinces us that it is not feasible if we demand an exact representation of the number of events. Nevertheless, there are some very efficient solutions that require only $\Theta(\log\log n)$ bits and guarantee sufficient accuracy for a wide range of applications. As examples, one can point most probabilistic counters – probabilistic structures well known in the literature since the seminal Morris’ paper [36] followed by its thorough mathematical analysis by Flajolet in [24]. They are used as building blocks in many space-efficient algorithms in the field of data mining or distributed data aggregation in networks or smart metering, just to mention a few ([2] or [6]).
@@
In this paper, we investigate probabilistic counters from the privacy-protection perspective. Our analysis is based on a differential privacy notion, which is commonly considered the only state-of-the-art approach. The differential privacy has the undeniable advantage of being mathematically rigorous and formally provable, contrary to previous anonymity-derived privacy definitions. This approach to privacy-preserving protocols can be used to give a formal guarantee for privacy that is resilient to any form of post-processing. For a survey about differential privacy properties, see [18] and references therein. Analysis of protocols based on differential privacy is usually technically complex, but by using this notion, we are immune to, e.g., linkage attacks (see, for example [37, 38]).
@@
The idea behind differential privacy is as follows: for two "neighbouring" scenarios that differ only in the participation of a single user, a differentially private mechanism should provide a response chosen from very similar distributions. Roughly speaking, differential privacy is described by two parameters: $\varepsilon$ – which controls a similarity of probabilities of common events – and $\delta$ – which is related to a probability of unusual events. The smaller the parameters, the better from the privacy point of view. In effect, judging by the output of the mechanism, one cannot say if a given individual (user) was taken into account for producing a given output. Intuitively, probabilistic counters should provide a high level of differential privacy since, statistically, many various numbers of events are "squeezed" into a small space of possible output results. In the case of one counter considered in our paper (MaxGeo) counter, one can find some similar, recent results about the privacy the algorithm offers. Nevertheless, the question about the value of the parameters of the potential differential privacy property remains open (see the discussion in Section 2) .
Moreover, when considering a small number of events $n$, an additional problem may occur, as it may be possible to distinguish that the number of events is different from n1n-1 or $n+1$ with a significant probability.
In our paper, we provide a very precise analysis of two well known probabilistic counters from the perspective of preserving privacy. It turned out that this task is surprisingly complex from the mathematical point of view.
@@
Our primary motivation is to find possibly accurate privacy parameters for the two most fundamental probabilistic counter protocols, namely the Morris Counter [36] and the MaxGeo Counter [47]. Note that the second one is used for yet another popular algorithm — HyperLogLog [22]. One may realise that these two counters are relatively old; however, they, together with their modifications, have been extensively used until these days. Morris Counter is often used in big data solutions, for instance, for measuring network’s capabilities [20]. The most crucial examples of refinements of the HyperLogLog algorithm are mentioned in Section 2.
@@
We claim that a high-precision analysis in the case of probabilistic counters is particularly important. This is because even a mechanism with very good privacy parameters can cause a serious privacy breach when used multiple times. That is, privacy loss/information leakage accumulates over multiple releases (see, e.g., [18]). Probabilistic counters in realistic scenarios may be used as fundamental primitives and subroutines in more complex protocols, since the differential privacy property is immune to post-processing.
@@
We also show that those two probabilistic counters can be used safely without any additional randomisation, even in very demanding settings. It is commonly known that no deterministic algorithm can provide non-trivial differential privacy. However, Probabilistic counters possess inherent randomness, achieving the desired privacy parameters. In other words, one can say that probabilistic counters are safe by design, and we do not need any additional privacy-orientated methods. In particular, what is most important, existing, working implementations do not need to be changed if we start demanding the provable privacy of a system.
@@
To the best of our knowledge, most of the results are new and deal with protocols not considered before in the context of privacy preservation. Some other (such as the MaxGeo counter) improves some previous results (e.g., [45]).
@@ 1.1 Paper structure and results
  • 4.
    We construct a privacy-preserving distributed survey protocol based on probabilistic counters in Section 5 and compare it with the Laplace method, which is considered as the actual state of the art of differentially private protocols and is not based on probabilistic counters.
The remainder of this paper is organised as follows. First, in Section 2 we mention work related to our paper and some popular examples of other probabilistic counters, which are not considered in this paper.
In Section 3, we recall the differential privacy definition. In Section 4 we describe probabilistic counters; further, we recall the definitions of both Morris and MaxGeo Counters. Moreover, we state Fact 1, a useful reformulation of the standard definition of differential privacy for probabilistic counters. In Section 5, we demonstrate how a probabilistic counter can be used to construct a data aggregation protocol in a very particular, yet natural scenario. Section 6 consists of formulations of our main results for both counters. Section 7 gives some ideas about the realisation of the scenario and the comparison of both counters. We also compare these solutions with the standard Laplace method (Section 7). For convenience of the reader, we provide the proofs in Section 8, since they are rather technical. For more clarity, some proofs and lemmas are moved to A. Finally, in Section 9 we present conclusions and future work propositions.
@@ 2 Previous and Related Work
In our paper, we focus on the inherent privacy guarantees of some probabilistic structures defined as differential privacy. The idea of differential privacy has been introduced for the first time in [16]; however, its precise formulation in the widely used form appeared for the first time in [19]. There is a long list of papers concerning differential privacy, e.g. [14, 15, 17], to mention a few.
Most of these papers focus on a centralised (global) model, namely, a database with a trusted party holding it. See that in our paper, despite the distributed setting, we have the same (non-local) trust model. In particular, we assume an existence of a curator that is entitled to gather and see all participants’ data in the clear and release the computed data to a wider (possibly untrusted) audience. Comprehensive information on differential privacy can be found in [18].
@@
The idea of probabilistic counters, along with the well-known Morris Counter was presented in the seminal paper [36]. The aim was to construct a very small data structure to represent a large set of events of some kind. In our paper, we focus on the Morris Counter analysed in detail in [24]. The second structure discussed in our paper is MaxGeo Counter, introduced and analysed in [47]. More detailed and precise analysis can be found in [21]. The most important application of MaxGeo Counter can be found in [22], where the authors propose the well-known HyperLogLog algorithm. Its practical applications are widely described in [28]. There are several widely used improvements of the HyperLogLog algorithm: HyperLogLog+ [28], Streaming HyperLogLog with sketches based on historical inverse probability [8] or martingal estimator [48] or empirically adjusted HyperBitBit (proposed by R. Sedgewick [44]). The main goal of these adjustments is a reduction of the memory requirements (see, e.g., [52] or [49]). For instance, some of the above solutions are used in database systems for query’ optimisation or for document classification purposes. Moreover, the MaxGeo counter was used in [41], for an adjustment of the ANF tool, developed for data mining from extensive graphs, which enables it to answer many different questions based on some neighbourhood function defined on the graph.
@@
Unsurprisingly, one of the main applications of the approximate counter is to compute the size of a database or its specific subset. A set of such applications can be found in [23]. In [50], the authors use Morris Counter for online, probabilistic, and space-efficient counting over streams of fixed, finite length. The authors of [7] proposed an application of a Morris Counter system for flash memory devices. Another application, presented in [10], is a revisit of Morris Counter designed for binary floating-point numbers. In [27], Morris Counter is used in a well-known problem of counting frequency moments of long data streams. The authors of [12] focused on making probabilistic counters scalable and accurate in concurrent settings. The paper on probabilistic counters in hardware can be found in [43]. A slightly modified version of Morris Counter called Morris+ was recently introduced in [39] with the proof of its optimality in terms of accuracy–memory trade-off.
@@
In random graphs theory, Morris Counter is usually connected to greedy structures. For instance, in an arrangement of a randomly labelled graph in Gilbert model $G(n,p)$, it is possible to construct a greedy stable set $S_{n}$, which size has the same distribution as Morris Counter $M_{n}$ of the base $a=(1-p)^{-1}$ (see, e.g., [25] or [4] for fundamentals of random graph theory).
@@
There are many other birth processes that are quite similar to the Morris Counter, which are applicable in a variety of disciplines like biology, physics, or the theory of random graphs. Short descriptions of such examples can be found in [9].
When talking about probabilistic counters, it is worth mentioning Bloom filter [3], which are space-efficient probabilistic data structures that are the representations of sets. There exists a probabilistic counter that approximates the number of elements represented by the given Bloom filter [46].
Other common examples of probabilistic counters are $F_{p}$ counters [1, 31], which approximate the $p$-th moments of frequencies of occurrences of different elements in the database.
@@
Let us also mention a paper [35] in which one can find numerous applications of similar constructions to create pseudorandom sketches in Big Data algorithms.
@@
Some probabilistic counters and similar structures were previously considered in terms of privacy preservation. We mention only the papers strictly related to the algorithms discussed in our paper (i.e., Morris Counter and MaxGeo).
The authors of [11] show that in the scenario of using different types of probabilistic counters for set cardinality estimation with the Adversary being able to extract the intermediate values of the counter, privacy is not preserved. Note that in this paper, we perform data aggregation instead of cardinality estimation. Moreover, we assume the Adversary is not able to extract any intermediate values from the counter. That is, we consider a global model, while the result from [11] assumes the settings closer to the classic local model [18].
One of the main results of this submission is a careful and tight analysis of Morris Counter from the context of preserving privacy. To the best of our knowledge, such analises have not been provided so far. Our second contribution is an analogous analysis of the MaxGeo.
There are a few very recent papers presenting privacy-preserving protocols that use the Flajolet–Martin sketch as a building block. In [23], so called Flajolet—Martin sketch was introduced. In [45] authors consider a general concept of a probabilistic counter, based on several MaxGeo counters and its differential privacy. Nevertheless, they incorrectly call it a generalisation of the Flajolet–Martin sketch (they probably confuse the Flajolet–Martin sketch with LogLog sketch).
@@
  • 1.
    In [45] the authors consider, among others, a sketch that can be seen as a particular application of the MaxGeo counter. They introduce its differentially private version via trick (adding artificial utilities) and provide its accuracy when used to count the number of elements in multisets.
    Accidentally, a proof of the basic theorem from [45] uses an incorrect argument (inappropriate utilisation of Hoeffding’s inequality), so it is difficult to compare the results precisely. Nevertheless, the overlap of results between our paper and [45] is only partial.
  • 2.
    In [5], the authors consider the LogLog sketch as a subroutine. After a careful analysis, they show that it is asymptotically $(\varepsilon,\delta=\mathrm{negl}(\lambda))$-DP (with respect to the numbers of different elements), when the number of elements counted by the mechanism is at least $8K\lambda\max(\frac{1}{\varepsilon},1)$, where $K$ is some accuracy parameter, $\lambda$ is some security parameter and $\mathrm{negl}(x)$ is some negligible function of argument $x$ (Theorem 4.2 in [5]). Nevertheless, the analysis does not explain how to choose parameters $K$ and $\lambda$ in order to obtain $(\varepsilon,\delta)$-DP for a given $\varepsilon$ and $\delta$ parameters. Moreover, a consideration of asymptotic behaviour (with respect to the number of unique elements $n$) is not relevant when the hash function restricts the possible result to the size bounded by its domain. Our analysis of the MaxGeo counter provides an exact (non-asymptotic) dependence between $n$ and the parameters $\varepsilon$ and $\delta$.
We also mention that some other pseudorandom structures have been analysed from the perspective of differential privacy. For example, in [32], the authors considered Bloom filters as a means for constructing privacy-preserving aggregation protocol.

### v7

# Probabilistic Counters for Privacy Preserving Data Aggregationthanks: Supported by Polish National Science Center grant number UMO-2018/29/B/ST6/02969UMO-2018/29/B/ST6/02969

@@ Abstract
We demonstrate that probabilistic counters can be used as a privacy protection mechanism without extra randomisation. That is, the inherent randomisation of the protocol is sufficient to protect privacy, even if the probabilistic counter is used multiple times. In particular, we present a specific privacy-preserving data aggregation protocol based on Morris Counter and MaxGeo Counter. Some of the results presented are devoted to counters that have not been investigated so far from the perspective of privacy protection. Another part is an improvement of the previous results. We show how our results can be used to perform distributed surveys and compare the properties of counter-based solutions and a standard Laplace method.
@@
probabilistic counter, Morris counter, differential privacy
@@
vol. 28:220261410.46298/dmtcs.116142023-07-20; 2023-07-20; 2026-02-122026-02-12
@@ 1 Introduction
Since Big Data related topics have been widely developed in recent years, solutions that focus on saving memory resources have become very popular. We would like to consider a standard example of such space-efficient mechanisms, namely probabilistic counters, which are used to represent the cardinality of dynamically counted events. More precisely, we would like to indicate the occurrence of $n$ events using a very small (significantly less than $\log n$) number of bits. We assume that $n$ is unknown in advance and may change. Clearly, a simple information-theoretic argument convinces us that it is not feasible if we demand an exact representation of the number of events. Nevertheless, there are some very efficient solutions that require only $\Theta(\log\log n)$ bits and guarantee sufficient accuracy for a wide range of applications. As examples, one can point most probabilistic counters – probabilistic structures well known in the literature since the seminal Morris’ paper Morris (1978) followed by its thorough mathematical analysis by Flajolet in Flajolet (1985). They are used as building blocks in many space-efficient algorithms in the field of data mining or distributed data aggregation in networks or smart metering, just to mention a few (Baquero et al. (2009) or Cichoń and Gotfryd (2018)).
@@
In this paper, we investigate probabilistic counters from the privacy-protection perspective. Our analysis is based on a differential privacy notion, which is commonly considered the only state-of-the-art approach. The differential privacy has the undeniable advantage of being mathematically rigorous and formally provable, contrary to previous anonymity-derived privacy definitions. This approach to privacy-preserving protocols can be used to give a formal guarantee for privacy that is resilient to any form of post-processing. For a survey about differential privacy properties, see Dwork and Roth (2014) and references therein. Analysis of protocols based on differential privacy is usually technically complex, but by using this notion, we are immune to, e.g., linkage attacks (see, for example Narayanan and Shmatikov (2009, 2010)).
@@
The idea behind differential privacy is as follows: for two ”neighbouring” scenarios that differ only in the participation of a single user, a differentially private mechanism should provide a response chosen from very similar distributions. Roughly speaking, differential privacy is described by two parameters: $\varepsilon$ – which controls a similarity of probabilities of common events – and $\delta$ – which is related to a probability of unusual events. The smaller the parameters, the better from the privacy point of view. In effect, judging by the output of the mechanism, one cannot say if a given individual (user) was taken into account for producing a given output. Intuitively, probabilistic counters should provide a high level of differential privacy since, statistically, many various numbers of events are ”squeezed” into a small space of possible output results. In the case of one counter considered in our paper (MaxGeo) counter, one can find some similar, recent results about the privacy the algorithm offers. Nevertheless, the question about the value of the parameters of the potential differential privacy property remains open (see the discussion in Section 2). Moreover, when considering a small number of events $n$, an additional problem may occur, as it may be possible to distinguish that the number of events is different from n1n-1 or $n+1$ with a significant probability.
@@
In our paper, we provide a very precise analysis of two well known probabilistic counters from the perspective of preserving privacy. It turned out that this task is surprisingly complex from the mathematical point of view. Our primary motivation is to find possibly accurate privacy parameters for the two most fundamental probabilistic counter protocols, namely the Morris Counter Morris (1978) and the MaxGeo Counter Szpankowski and Rego (1990). Note that the second one is used for yet another popular algorithm — HyperLogLog Flajolet et al. (2007). One may realise that these two counters are relatively old; however, they, together with their modifications, have been extensively used until these days. Morris Counter is often used in big data solutions, for instance, to measure network’s capabilities Einziger et al. (2018). The most crucial examples of refinements of the HyperLogLog algorithm are mentioned in Section 2.
@@
We claim that a high-precision analysis in the case of probabilistic counters is particularly important. This is because even a mechanism with very good privacy parameters can cause a serious privacy breach when used multiple times. That is, privacy loss/information leakage accumulates over multiple releases (see, e.g., Dwork and Roth (2014)). Probabilistic counters in realistic scenarios may be used as fundamental primitives and subroutines in more complex protocols, since the differential privacy property is immune to post-processing.
@@
We also show that those two probabilistic counters can be used safely without any additional randomisation, even in very demanding settings. It is commonly known that no deterministic algorithm can provide non-trivial differential privacy. However, Probabilistic counters have inherent randomness, achieving the desired privacy parameters. In other words, one can say that probabilistic counters are safe by design, and we do not need any additional privacy-orientated methods. In particular, what is most important, existing, working implementations do not need to be changed if we start demanding the provable privacy of a system.
@@
To the best of our knowledge, most of the results are new and deal with protocols not considered before in the context of privacy preservation. Some other (such as the MaxGeo counter) improve some previous results (e.g. Smith et al. (2020)).
@@ 1.1 Paper structure and results
  • We construct a distributed survey protocol to preserve privacy based on probabilistic counters in Section 5 and compare it with the Laplace method, which is considered as the actual state of the art of differentially private protocols and is not based on probabilistic counters.
The remainder of this paper is organised as follows. First, in Section 2 we mention work related to our paper and some popular examples of other probabilistic counters, which are not considered in this paper. In Section 3, we recall the differential definition of privacy. In Section 4 we describe probabilistic counters; further, we recall the definitions of both Morris and MaxGeo Counters. Moreover, we state Fact 1, a useful reformulation of the standard definition of differential privacy for probabilistic counters. In Section 5, we demonstrate how a probabilistic counter can be used to construct a data aggregation protocol in a very particular, yet natural scenario. Section 6 consists of formulations of our main results for both counters. Section 7 gives some ideas about the realisation of the scenario and the comparison of both counters. We also compare these solutions with the standard Laplace method (Section 7). For convenience of the reader, we provide the proofs in Section 8, as they are rather technical. For more clarity, some proofs and lemmas are moved to A. Finally, in Section 9 we present conclusions and future work propositions.
@@ 2 Previous and Related Work
In our paper, we focus on the inherent privacy guarantees of some probabilistic structures defined as differential privacy. The idea of differential privacy has been introduced for the first time in Dwork et al. (2006b); however, its precise formulation in the widely used form appeared for the first time in Dwork (2006). There is a long list of papers concerning differential privacy, e.g. Dwork et al. (2006a); Dwork and Lei (2009); Dwork et al. (2010), to mention a few.
Most of these papers focus on a centralised (global) model, namely, a database with a trusted party holding it. See that in our paper, despite the distributed setting, we have the same (non-local) trust model. In particular, we assume an existence of a curator that is entitled to gather and see all participants’ data in the clear and release the computed data to a wider (possibly untrusted) audience. Comprehensive information on differential privacy can be found in Dwork and Roth (2014).
@@
The idea of probabilistic counters, along with the well-known Morris Counter was presented in the seminal paper Morris (1978). The aim was to construct a very small data structure to represent a large set of events of some kind. In our paper, we focus on the Morris Counter analysed in detail in Flajolet (1985). The second structure discussed in our paper is MaxGeo Counter, introduced and analysed in Szpankowski and Rego (1990). More detailed and precise analysis can be found in Eisenberg (2008). The most important application of MaxGeo Counter can be found in Flajolet et al. (2007), where the authors propose the well-known HyperLogLog algorithm. Its practical applications are widely described in Heule et al. (2013). There are several widely used improvements of the HyperLogLog algorithm: HyperLogLog+ Heule et al. (2013), Streaming HyperLogLog with sketches based on historical inverse probability Cohen (2015) or martingal estimator Ting (2014) or empirically adjusted HyperBitBit (proposed by R. Sedgewick Sedgewick (2018)). The main goal of these adjustments is to reduce the memory requirements (see, e.g., Yun et al. (2015) or Ting (2019)). For instance, some of the above solutions are used in database systems for query’ optimisation or for document classification purposes. Moreover, the MaxGeo counter was used in Palmer et al. (2002), for an adjustment of the ANF tool, developed for data mining from extensive graphs, which enables it to answer many different questions based on some neighbourhood function defined on the graph.
@@
Unsurprisingly, one of the main applications of the approximate counter is to compute the size of a database or its specific subset. A set of such applications can be found in Flajolet and Martin (1985). In Van Durme and Lall (2009), the authors use Morris Counter for online, probabilistic, and space-efficient counting over streams of fixed, finite length. The authors of Cichon and Macyna (2011) proposed an application of a Morris Counter system for flash memory devices. Another application, presented in Csűrös (2010), is a revisit of Morris Counter designed for binary floating-point numbers. In Gronemeier and Sauerhoff (2009), Morris Counter is used in a well-known problem of counting frequency moments of long data streams. The authors of Dice et al. (2013) focused on making probabilistic counters scalable and accurate in concurrent settings. The paper on probabilistic counters in hardware can be found in Riley and Zilles (2006). A slightly modified version of Morris Counter called Morris+ was recently introduced in Nelson and Yu (2022) with the proof of its optimality in terms of accuracy–memory trade-off.
@@
In random graphs theory, Morris Counter is usually connected to greedy structures. For instance, in an arrangement of a randomly labelled graph in Gilbert model $G(n,p)$, it is possible to construct a greedy stable set $S_{n}$, which size has the same distribution as the Morris Counter $M_{n}$ of the base $a=(1-p)^{-1}$ (see, e.g., Frieze and Karonski (2015) or Bollobás (1998) for the fundamentals of random graph theory).
@@
There are many other birth processes that are quite similar to the Morris Counter, which are applicable in a variety of disciplines like biology, physics, or the theory of random graphs. Short descriptions of such examples can be found in Crippa and Simon (1997). When talking about probabilistic counters, it is worth mentioning the Bloom filter Bloom (1970), which are space-efficient probabilistic data structures that are representations of sets. There exists a probabilistic counter that approximates the number of elements represented by the given Bloom filter Swamidass and Baldi (2007).
@@
Other common examples of probabilistic counters are $F_{p}$ counters Alon et al. (1999); Indyk (2000), which approximate the $p$-th moments of frequencies of occurrences of different elements in the database. Let us also mention a paper Mishra and Sandler (2006) in which one can find numerous applications of similar constructions to create pseudorandom sketches in Big Data algorithms.
@@
Some probabilistic counters and similar structures were previously considered in terms of privacy preservation. We mention only the papers strictly related to the algorithms discussed in our paper (i.e., Morris Counter and MaxGeo). The authors of Desfontaines et al. (2019) show that in the scenario of using different types of probabilistic counters for set cardinality estimation with the Adversary being able to extract the intermediate values of the counter, privacy is not preserved. Note that in this paper, we perform data aggregation instead of cardinality estimation. Moreover, we assume the Adversary is not able to extract any intermediate values from the counter. That is, we consider a global model, while the result from Desfontaines et al. (2019) assumes the settings closer to the classic local model Dwork and Roth (2014).
One of the main results of this submission is a careful and tight analysis of Morris Counter from the context of preserving privacy. To the best of our knowledge, such analises have not been provided so far. Our second contribution is an analogous analysis of the MaxGeo. There are a few very recent papers presenting privacy-preserving protocols that use the Flajolet–Martin sketch as a building block. In Flajolet and Martin (1985), so called Flajolet—Martin sketch was introduced. In Smith et al. (2020) authors consider a general concept of a probabilistic counter, based on several MaxGeo counters and its differential privacy. Nevertheless, they incorrectly call it a generalisation of the Flajolet–Martin sketch (they probably confuse the Flajolet–Martin sketch with LogLog sketch).
@@
In Smith et al. (2020) the authors consider, among others, a sketch that can be seen as a particular application of the MaxGeo counter. They introduce its differentially private version via trick (adding artificial utilities) and provide its accuracy when used to count the number of elements in multisets. Accidentally, a proof of the basic theorem from Smith et al. (2020) uses an incorrect argument (inappropriate utilisation of Hoeffding’s inequality), so it is difficult to compare the results precisely. Nevertheless, the overlap of results between our paper and Smith et al. (2020) is only partial.
In Choi et al. (2020), the authors consider the LogLog sketch as a subroutine. After a careful analysis, they show that it is asymptotically $(\varepsilon,\delta=\mathrm{negl}(\lambda))$-DP (with respect to the numbers of different elements), when the number of elements counted by the mechanism is at least $8K\lambda\max(\frac{1}{\varepsilon},1)$, where $K$ is some accuracy parameter, $\lambda$ is some security parameter and $\mathrm{negl}(x)$ is some negligible function of argument $x$ (Theorem 4.2 in Choi et al. (2020)). Nevertheless, the analysis does not explain how to choose parameters $K$ and $\lambda$ in order to obtain $(\varepsilon,\delta)$-DP for a given $\varepsilon$ and $\delta$ parameters. Moreover, a consideration of asymptotic behaviour (with respect to the number of unique elements $n$) is not relevant when the hash function restricts the possible result to the size bounded by its domain. Our analysis of the MaxGeo counter provides an exact (non-asymptotic) dependence between $n$ and the parameters $\varepsilon$ and $\delta$.
We also mention that some other pseudorandom structures have been analysed from the perspective of differential privacy. For example, in Klonowski and Piotrowska (2018), the authors considered Bloom filters as a means of constructing a privacy-preserving aggregation protocol.